11/5 divided by {21/3-(5+2-3)}-3/12
 

Grade:7

2 Answers

Prateek Pravanjan
8 years ago
 
11/5                    =11/5           (use BracketOfDivisionMultiplicationAdditionSubtraction rule)
[21/3-(5+2-3)]-3/12      [7-4-1/4]
                           
                            =11/5           
                              [3-1/4]
                            =11/5       {take 12 as common in the denominator}
                              [12/4-1/4]
                            =11/5
                              11/4
                            =11X4  (take 5 to the numerator and 4 to the denominator)
                              11X5   
                            =4/5 is the answer(it can also be written as 0.8)
                             HOPE YOU GET IT
A M S ARUN KRISHNA
8 years ago
11/5 divided by {21/3-(5+2-3)}-3/12
using BODMAS rule find {21/3-(5+2-3)}-3/12
BODMAS – Brackets Off Divide Multiply Add Subtract
= {21/3-(5+2-3)}-3/12
= 7 - 4 – ¼
= 3 – 1/4
= 11/4
        11/5            11        4
=  ------------  =  ------- × -------  =  4/5
        11/4             5         11
please approve my anwer